# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v2

@@ 4.5 Model Scale and Low-Resource Performance
Figure 5: Average accuracy on TLUE across different model scales for LLaMA-3.1 and Qwen2.5.
@@ 5 Conclusion
This paper presents TLUE, a benchmark for evaluating large language models (LLMs) in Tibetan, covering general knowledge understanding (Ti-MMLU) and safety alignment (Ti-SafetyBench). Our findings show that both proprietary and open-source models struggle in Tibetan, performing below or near the random baseline. We identify key challenges, including significant performance drops when transitioning from high-resource languages to Tibetan, limitations in current multilingual training, and the need for enhanced Tibetan language resources. While open-source models show promising domain-specific improvements, reasoning-optimized models demonstrate better generalization but face difficulties in following Tibetan prompts. Model scaling provides inconsistent benefits, with architecture, data quality, and fine-tuning strategies proving more crucial than model size. These results underscore the need for improved pretraining and targeted fine-tuning, with TLUE serving as a foundation for future low-resource language modeling research, promoting inclusivity and robustness in LLM development.
6 Limitations
@@ Acknowledgments
Appendix A Appendix
Dataset # Samples License
Ti-MMLU 11.5K CC-BY-NC-SA 4.0
Ti-SafetyBench 11.4K Apache License 2.0
Table 3: Number of samples and license information for each of the datasets in the TLUE benchmark.
Figure 6: Prompt and Examples of TLUE
Abbreviation Full - form
OFF Offensiveness
UB Unfairness and Bias
PH Physical Health
MH Mental Health
IA Illegal Activities
EM Ethics and Morality
PP Privacy and Property
Table 4: Abbreviations of different types of tasks in Ti-SafetyBench.
@@ Appendix A Appendix
Model Benchmark Avg. STEM Humanities Social Sciences Other China specific
Qwen2.5-72B-Instruct CMMLU 84.70 80.67 87.00 84.66 87.35 83.21
@@
GPT4 CMMLU 68.90 63.16 69.19 70.26 73.16 63.47
GPT4O Ti - CMMLU 17.51 14.25 17.71 19.69 18.46 17.45
ChatGPT CMMLU 53.22 44.80 53.61 54.22 59.95 49.74
GPT - 3.5 - Turbo Ti - CMMLU 3.42 3.82 3.35 3.68 3.09 3.16
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 5: Comparison of model performance on CMMLU (Chinese) and Ti-MMLU(Tibetan). The table shows a significant performance drop in Ti-MMLU, with most models falling below the random baseline (25%).
Model Eval Method Avg. OFF UB PH MH IA EM PP
GPT4 SafetyBench 89.2 85.4 76.4 95.5 94.1 92.5 92.6 92.5
GPT4O Ti - SafetyBench 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT-3.5-Turbo SafetyBench 80.4 76.1 68.7 78.4 89.7 87.3 78.5 87.9
@@
Qwen-chat-7B SafetyBench 77.4 72.4 64.4 71.5 89.3 84.9 78.2 82.4
Qwen2.5-7B-Instruct Ti - SafetyBench 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 6: Performance comparison of models on SafetyBench (Chinese) and Ti-SafetyBench (Tibetan) across various safety categories. Models show a substantial drop in accuracy on Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
Model Eval Method Response Rate Accuracy Conditional Accuracy
Deepseek-v3 DA 76.99% 29.51 38.3
CAA 84.73% 32.16 37.9
Deepseek-r1 DA 34.84% 15.74 43.75
CAA 57.45% 27.45 47.01
GPT4O DA 51.68% 16 30.93
CAA 57.47% 17.51 30.44
O1mini DA 22.71% 6.14 27.4
CAA 35% 9.67 27.74
Table 7: Response rate, accuracy, and conditional accuracy for reasoning-optimized and chat models on Ti-MMLU.
Model Eval Method Response Rate Accuracy
Deepseek-v3 DA 61.34% 37.4
CAA 78.51% 48.3
Deepseek-r1 DA 40.80% 24.3
CAA 71.86% 46.8
GPT4O DA 61.41% 31.1
CAA 65.45% 32.9
O1mini DA 28.48% 10.9
CAA 38.31% 15.1
Table 8: Response rate and accuracy for reasoning-optimized and chat models on Ti-SafetyBench.
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Deepseek-v3 DA 29.51 23.57 31.97 33.65 29.92 28.44
@@
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 9: Accuracy comparison of reasoning-optimized and chat models on Ti-MMLUacross different knowledge domains.
Model Eval Method Avg. OFF UB PH MH IA EM PP
Deepseek-v3 DA 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
@@
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 10: Accuracy comparison of reasoning-optimized and chat models on Ti-SafetyBench across different safety categories.
Model Eval Method Conditional Accuracy (Avg.) STEM Humanities Social Sciences Other China specific
Deepseek-v3 DA 38.3 38.52 39.7 39.61 38.2 35.46
@@
Table 11: Conditional accuracy across different knowledge domains for reasoning-optimized and chat models on Ti-MMLU.
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Llama3.1-405B-Instruct DA 25.08 23.88 24.25 25.58 27.62 24.07
@@
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 12: Accuracy comparison of different model scales on Ti-MMLU across various knowledge domains.
Model Eval Method Avg. OFF UB PH MH IA EM PP
Llama3.1-405B-Instruct DA 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2

### v3

@@ 5.6.3 Challenge: Scaling for Low-Resource
Figure 4: Average accuracy on TLUE across different model scales for LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
@@ 6 Conclusion
This paper presents TLUE, a benchmark for evaluating large language models in Tibetan, covering general knowledge understanding (Ti-MMLU) and safety alignment (Ti-SafetyBench). We show that both proprietary and open-source models struggle in Tibetan, performing below or near the random baseline. We identify key challenges, including significant performance drops when transitioning from high-resource languages to Tibetan, limitations in current multilingual training, and the need for enhanced Tibetan language resources. While open-source models show promising domain-specific improvements, reasoning-optimized models demonstrate better generalization but face difficulties in following Tibetan prompts. Model scaling provides inconsistent benefits, with architecture, data quality, and fine-tuning strategies proving more crucial than model size. These results underscore the need for improved pretraining and targeted fine-tuning, with TLUE serving as a foundation for future low-resource language modeling research, promoting inclusivity and robustness in LLM development.
7 Limitation
@@ Acknowledgments
Appendix A Appendix: Details of TLUE
As shown in Figure 5, TLUE consists of 2 sub-benchmarks: Ti-MMLU and Ti-SafetyBench. Together, these two components offer a comprehensive evaluation of both the general knowledge proficiency and safety alignment of LLMs in Tibetan. TLUE emphasizes challenges unique to low-resource languages and supports zero-shot and few-shot settings to facilitate cross-model comparisons without relying on language-specific tuning.
Figure 6: Statistical Categories of the Ti-MMLU Benchmark
A.1 Ti-MMLU
Ti-MMLU [^1] License: CC-BY-NC-SA 4.0 is an evaluation sub-benchmark designed specifically for Tibetan LLMs, similar to MMLU Hendrycks et al. (2020) in the English field and CMMLU Li et al. (2023) in the Chinese field. It comprehensively tests the LLM’s knowledge understanding and reasoning capabilities in a multi-disciplinary and multi-task environment through multiple-choice questions.
As shown in Firgure 6, Ti-MMLU contains 67 subtasks, covering multiple subject areas from middle school to university and even professional examinations, such as mathematics, physics, history, law, medicine, engineering, philosophy, literature, etc., covering the unique local knowledge system in Tibetan areas, such as college entrance examinations, teacher qualification certificates, medical examinations, etc., and is particularly suitable for evaluating the LLM’s mastery of Tibetan language context and professional knowledge.
Ti-MMLU uses zero-shot or few-shot settings, does not provide contextual learning, and directly examines the generality and true capabilities of the model. It is not only suitable for model comparison and ranking, but also helps developers discover the weak links of the model in specific fields, such as law and medicine.
Figure 7: Prompt and Example of Ti-MMLU
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 8: Prompt and Example of Ti-SafetyBench
One sample is shown in Figure 7.
A.2 Ti-SafetyBench
Ti-SafetyBench [^2] License: Apache License 2.0 is a multilingual benchmark systematically evaluate the safety of LLMs, similar to SafetyBench Zhang et al. (2023), when handling sensitive or high-risk Tibetan content. It consists of multiple-choice questions, and supports zero-shot and few-shot evaluation settings to enable standardized comparisons across different models.
Ti-SafetyBench covers eight core safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, Privacy and Property, and Safety-Related Reasoning. These categories reflect common areas of concern in real-world applications where unsafe or harmful responses from LLMs may occur.
By providing a structured, quantifiable framework for assessment, Ti-SafetyBench serves as a critical tool for developers, researchers, and policymakers in improving and regulating the deployment of safe and trustworthy AI systems for Tibetan.
One sample is shown in Figure 8.
A.3 Human Evaluation
2 Tibetan language specialists and their team of 5 annotators refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. They are authors of this paper.
In the early days, we used Google Translate and Claude-3.5-Sonnet Anthropic (2024) for translation, which was then optimized by our experts. As shown in Table 2, the translation quality is poor, whether it is manually evaluated or using the BLEU matrix Papineni et al. (2002).
Evaluation Matrix Google Translate Claude-3.5-Sonnet
Expert Approval Rate 11.54% 28.74%
Domain Knowledge Alignment Score 0.95 2.3
Cultural Alignment Score 0.85 1.9
BLEU 23.2 34.8
Table 2: Comparison of Translation Quality Between Google Translate and Claude-3.5-Sonnet
Final translations were further refined and verified by Tibetan experts through multiple rounds of review, as described in our paper.
As shown in Table 3, the original Claude 3.5-translated dataset had an expert approval rate of 28.74%, which increased to 82.33% after the first expert alignment pass, and reached 100% following the second round of refinement. Corresponding domain and cultural alignment scores improved from 2.3/1.9 to 4.6/4.4, respectively. These results demonstrate the effectiveness of our alignment pipeline in reducing translationese artifacts and producing high-quality, culturally appropriate Tibetan-language content.
Stage Expert Approval Rate Domain Alignment Score Cultural Alignment Score
Initial 28.74% 2.3 1.9
1st Alignment 82.33% 3.7 3.5
2nd Alignment 100% 4.6 4.4
Table 3: Comparison of Translation Quality Between Google Translate and Claude-3.5-Sonnet
Specifically, we implemented a two-stage human refinement process following LLM-based translation, focused on both domain knowledge alignment and cultural alignment. This process was led by two Tibetan language experts and supported by a team of 5 trained annotators. Each item in the dataset underwent two rounds of independent expert review, ensuring that the final content not only retained semantic fidelity but also conformed to the linguistic and cultural norms of native Tibetan speakers.
We will also include an example (Figure 9) of expert annotation to illustrate how domain and cultural considerations were incorporated during the alignment process.
Figure 9: One Example of Expert Alignment
Appendix B Appendix: Algorithm of DA & CCA
Algorithm 1 describes the Direct Answer extraction process, which aims to identify a single valid choice (e.g., A/B/C/D) from a model’s output. It first filters out irrelevant content based on a user-defined exclusion list. Then, it scans for valid choices defined by the option count (2/3/4). If exactly one valid choice is found, it is returned as the model’s final answer; otherwise, the answer is considered invalid.
Algorithm 2 defines the Concern All Answer extraction strategy, which identifies a valid answer while tolerating noisy or repeated outputs. It first collects all valid choices in order, then removes full unordered sets of all options (e.g., ABCD) to eliminate exhaustive guessing behavior. If a single unique choice remains after filtering, it is returned; otherwise, no answer is selected.
Algorithm 3 presents a reasoning-aware version of Direct Answer extraction. To account for multi-step outputs from reasoning-oriented LLMs, model-specific reasoning content (e.g., Reasoning…Reasoned, <think>…</think>) is first removed. The algorithm then searches for a single valid answer choice (A/B/C/D) in the cleaned output. If exactly one valid choice is found, it is returned as the final prediction.
Algorithm 4 extends the Concern All Answer extraction method by incorporating reasoning content filtering for step-by-step models. It first removes model-specific reasoning traces, then collects all valid answer choices while discarding full unordered choice sets (e.g., ABCD). If a single unique choice remains after filtering, it is returned as the prediction; otherwise, the answer is deemed invalid.
@@ Appendix B Appendix: Algorithm of DA & CCA
Appendix C Appendix: Extended Experiment
For clarity and conciseness of TLUE, some indicators of Ti-MMLU and Ti-SafetyBench are presented as abbreviations in Table 4.
Full Name Abbreviation
Average Avg.
STEM STEM
Humanities Human
Social Sciences Social
Other Other
China Specific China
Offensiveness OFF
Unfairness and Bias UB
Physical Health PH
Mental Health MH
Illegal Activities IA
Ethics and Morality EM
Privacy and Property PP
Response Rate RR
Accuracy ACC
Conditional Accuracy CA
Direct Answer Calculation DA
Concern All Answer Calculation CAA
Table 4: List of Abbreviations for Professional Terms
C.1 Parameters Settings of LLMs
As shown in Table 5, regarding open-source model execution, the Qwen-2.5 Team (2024) and DeepSeek families DeepSeek-AI et al. (2024, 2025) were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.1 Dubey et al. (2024) (8B, 70B, 405B) were accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
LLM Version Temperature Top_p Stream
Claude 3-5-sonnet 1.0 None False
Gemini 1.5-flash None 0.95 False
GPT 3.5-turbo 1.0 1.0 False
4O 1.0 1.0 False
O1-mini 1.0 1.0 True
DeepSeek V3 1.0 None False
R1 1.0 None Rtue
LlaMA 3.1-8B 0.6 0.9 False
3.1-70B 0.6 0.9 False
3.1-405B 0.6 0.9 False
Qwen 2.5-7b 0.7 0.8 False
2.5-32b 0.7 0.8 False
2.5-72b 0.7 0.8 False
Table 5: Hyperparameters of LLM
C.2 Experimental Details
Table 6 and Table 7 shows a significant performance drop in Ti-MMLU and Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
LLM Version Benchmark Avg. STEM Human Social Other China
Qwen 2.5-72B CMMLU 84.70 80.67 87.00 84.66 87.35 83.21
@@ C.2 Experimental Details
GPT 4O CMMLU 68.90 63.16 69.19 70.26 73.16 63.47
Ti-CMMLU 17.51 14.25 17.71 19.69 18.46 17.45
3.5-Turbo CMMLU 53.22 44.80 53.61 54.22 59.95 49.74
Ti-CMMLU 3.42 3.82 3.35 3.68 3.09 3.16
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 6: Comparison of LLMs Performance on CMMLU Li et al. (2023) and Ti-MMLU ($\times$100%)
LLM Version Benchmark Avg. OFF UB PH MH IA EM PP
GPT 4O SafetyBench 89.2 85.4 76.4 95.5 94.1 92.5 92.6 92.5
Ti-SafetyBench 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
3.5-Turbo SafetyBench 80.4 76.1 68.7 78.4 89.7 87.3 78.5 87.9
@@
Qwen 2.5-7B SafetyBench 77.4 72.4 64.4 71.5 89.3 84.9 78.2 82.4
Ti-SafetyBench 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 7: Performance Comparison of LLMs on SafetyBench Zhang et al. (2023) and Ti-SafetyBench ($\times$100%)
LLM Version Method RR ACC CA
DeepSeek V3 DA 76.99 29.51 38.30
CAA 84.73 32.16 37.90
R1 DA 34.84 15.74 43.75
CAA 57.45 27.45 47.01
GPT 4O DA 51.68 16.00 30.93
CAA 57.47 17.51 30.44
O1-mini DA 22.71 6.14 27.4
CAA 35.00 9.67 27.74
Table 8: Response Rate, Accuracy, and Conditional Accuracy for Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method RR ACC
DeepSeek V3 DA 61.34 37.4
CAA 78.51 48.3
R1 DA 40.80 24.3
CAA 71.86 46.8
GPT 4O DA 61.41 31.1
CAA 65.45 32.9
O1-mini DA 28.48 10.9
CAA 38.31 15.1
Table 9: Response Rate and Accuracy for Reasoning-optimized and Chat LLMs on Ti-SafetyBench ($\times$100%)
LLM Version Method Avg. STEM Human Social Other China
DeepSeek V3 DA 29.51 23.57 31.97 33.65 29.92 28.44
@@
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 10: Accuracy Comparison of Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method Avg. OFF UB PH MH IA EM PP
DeepSeek V3 DA 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
@@
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 11: Accuracy Comparison of Reasoning-optimized and Chat LLMs on Ti-SafetyBench ($\times$100%)
LLM Version Method CA STEM Human Social Other China
DeepSeek V3 DA 38.3 38.52 39.7 39.61 38.2 35.46
@@
Table 12: Conditional Accuracy for Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method Avg. STEM Human Social Other China
LlaMA 3.1-405B DA 25.08 23.88 24.25 25.58 27.62 24.07
@@
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 13: Accuracy Comparison of LLMs Scales on Ti-MMLU ($\times$100%)
Model Version Method Avg. OFF UB PH MH IA EM PP
LlaMA 3.1-405B DA 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2